# On the conditioning of histopolationL. Bruni Bruno, S. Serra-Capizzano

### v1

@@ 2.3 Segments with uniform length: the class (C4)
The class (C4) shows that the way segments overlap is a rather weak information for the purpose of unisolvence. In fact, we are going to show that the class (C4) is unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$ is unisolvent whenever the nodes $\xi_{i}$, $i=1,\ldots,d$, are pairwise distinct, with no restriction on the measure of $s$. This can be seen as a consequence of the following general fact.
@@
Let $f:\mathbb{R}\to\mathbb{R}$ be a continuous function, and let $s$ be an interval. Define $s+\xi:=\{x+\xi:\ x\in s\}$. If
$\int_{s+\xi}f(x)\mathrm{d}x=0$
for each $x\in\mathbb{R}$, then $f(x)=0$ for each $x\in\mathbb{R}$.
Proof 2.9.
For ease, observe that the result can be restated in terms of translations by $\xi$ in place of $\xi$ and consider the characteristic function $\chi_{s-\xi}$ for the interval $s-\xi$. Thus, for each $\xi\in\mathbb{R}$, one has
$0=\int_{s-\xi}f(x)\mathrm{d}x=\int_{\mathbb{R}}\left(f\cdot\chi_{s-\xi}\right)(x)\mathrm{d}x=\left(f\ast\chi_{s-\xi}\right)(x),$
begin $\ast$ the usual convolution operator. By construction, one has that $f\cdot\chi_{s-\xi}(x)\in L^{1}(\mathbb{R})$, and we may consider the action of the Fourier transform $\mathcal{F}$ between integrable functions in $L^{1}$ and bounded functions in $L^{\infty}$, which gives
(6) $0=\mathcal{F}(0)=\mathcal{F}(f\ast\chi_{s-\xi})=\mathcal{F}(f)\cdot\mathcal{F}(\chi_{s-\xi}).$
By direct computation, setting $s=[\alpha,\beta]$, one obtains
$\left(\mathcal{F}(\chi_{s-\xi})\right)(t)=e^{i\pi t\left(2\xi-(\alpha+\beta)\right)}\frac{\sin\left(|s|t\right)}{t},$
continuously extended at $t=0$. The norm of this function is bounded by the measure of $s$, since $|\mathcal{F}(\chi_{s-\xi})|\leq|s-\xi|=|s|<+\infty$. Further, for each $\xi$ the function $\mathcal{F}(\chi_{s-\xi})$ vanishes on a discrete set. By (6), it must be $\mathcal{F}(f)=0$ almost everywhere, and the continuity of $f$ implies that $f=0$.
@@
Let $s$ be a segment, and define $s+\xi:=\{x+\xi:\ x\in s\}$. Consider $\left\{\xi_{1},\ldots,\xi_{d}\right\}$ pairwise distinct nodes and let $p_{d-1}\in\mathscr{P}_{d-1}(I)$. Then, if
@@
By the change of variable, this is a polynomial map (in the unknown $\xi$) of degree d1d-1. By hypothesis, this polynomial vanishes on the grid $\left\{\xi_{1},\ldots,\xi_{d}\right\}$, which consists of distinct nodes. Hence this polynomial vanishes on the whole real line $\mathbb{R}$. By Lemma 2.8, one gets $p_{d-1}(x)=0$ for each $x\in\mathbb{R}$.
@@
Proposition 2.10 also has a different proof. To this end, observe that the set of zeros of $p_{d-1}(x)$ is a discrete set, containing at most d1d-1 nodes. If it had more, then $p_{d-1}$ would be the zero polynomial. Call $\xi_{\max}$ the largest one, i.e. the node $\xi_{i}$ such that $\xi_{i}\geq\xi_{j}$ for any $i\neq j$. Then, the result of Lemma 2.8 can be deduced in the specific case of polynomials as there exists $\xi\in\mathbb{R}$ such that $\xi-|s|\geq\xi_{\max}$. Hence, the sign of $p_{d-1}$ is constant in the interior of $s+\xi$, and since $\int_{s+\xi}p_{d-1}(x)\mathrm{d}x=0$, one has $p_{d-1}(x)=0$ for each $x$ in the (non empty) interior of $s+\xi$. This implies that $p_{d-1}(x)=0$ for each $x\in\mathbb{R}$.
@@ 2.4 Extension to generic intervals
and $\varepsilon$ is chosen so that $i+\varepsilon<2$.
@@
The picture given in Theorems 2.5 and 2.13 and in Remarks 2.1 and 2.15 is even more general. In fact, if we consider sets $\{i\cdot\alpha_{i}^{(d)}=1,\ldots,d\}$, $\{i\cdot\beta_{i}^{(d)}=1,\ldots,d\}$, of natural numbers, in which for $d_{1}\neq d_{2}$ we have in general $\alpha_{i}^{(d_{1})}\neq\alpha_{i}^{(d_{2})}$ and $\beta_{i}^{(d_{1})}\neq\beta_{i}^{(d_{2})}$, $i\leq\min\{d_{1},d_{2}\}$, all the proofs work unchanged. This represents a further degree of generality in the analysis.
@@ 2.5 Some remarks on the independence of $c$ on $d$
We have so far assumed the existence of a “separation distance” $c>0$ such that $\max_{i}\left(\beta_{i}-\alpha_{i}\right)\geq c$. We asked such a constant to be independent of $d$. This is clearly motivated by the fact that, if $c$ depended on $d$, it could decay to zero fast enough to simplify the exponential term $2^{d-2}$. It is plain to observe that, if $c$ depends on $d$ and does not decay exponentially, the exponential trend of the conditioning of the monomial basis remains unchanged. An example is provided when considering equidistributed segments in I=[1,1]I=[-1,1].
Suppose, in complete contrast, that such a constant $c=\max_{i}(\beta_{i}-\alpha_{i})$ is equal to zero; this means that every segment $s_{i}$ collapses to a point $\xi_{i}$. Even if these nodes are distinct, the problem (1) becomes ill-posed, as the matrix (2) becomes the zero matrix. This situation can be resolved by placing the normalization factor $|s_{i}|$ in the histopolation conditions, as discuessed in [5]. An immediate continuity argument, together with the mean value theorem, shows that the Vandermonde matrix (2) converges (in an entry-wise sense) to the nodal Vandermonde matrix based upon the nodes $\{\xi_{1},\ldots,\xi_{d}\}$. The well-established nodal theory [2, 19] hence ensures that Theorem 2.13 holds even in this limit.
@@ 3 Well-conditioning: Chebyshev basis and segments
In stark contrast with Section 2, a very different situation appears when the Chebyshev polynomials of the second kind
$U_{j}(x)=\frac{\sin(j\arccos x)}{\sin(\arccos x)}$
@@
is usually adopted. The two families are related via differentiation, as $\mathrm{d}/\mathrm{d}x\left(T_{j}(x)\right)=jU_{j-1}(x)$. The Vandermonde matrix (2) assumes a very peculiar form with respect to the basis $U_{j}(x)$:
(7) $\left[V_{d}\right]_{i,j}$ $=\int_{\alpha_{i}}^{\beta_{i}}U_{j}(x)\mathrm{d}x=\frac{1}{j}\int_{\alpha_{i}}^{\beta_{i}}T_{j+1}^{\prime}(x)\mathrm{d}x$
@@
(8) $s_{i}=[\alpha_{i},\beta_{i}]=[\cos(\tau_{i}+\rho),\cos(\tau_{i}-\rho)],$
being $\tau_{i}$ the (pre-image of the) $i$-th Chebyshev node $\tau_{i}=\frac{(i+1/2)\pi}{d}$ for $i=0,\ldots,d-1$ and $\rho>0$ an arc length veryfing the existence conditions established in [4, Section 3.2]. Plugging the definition of the polynomial $U_{j}(x)$ and class (C2) segments into (7), one immediately obtains that the respective Vandermonde matrix is
(9) $\left[V_{d}\right]_{i,j}=\frac{2}{j}\sin(j\tau_{i})\sin(j\rho).$
@@
(10) $\left[V_{d}^{\top}V_{d}\right]_{i,i}=\frac{2d}{i^{2}}\sin^{2}(i\rho).$
Proof 3.2.
@@
$\left[V_{d}^{\top}V_{d}\right]_{i,j}$ $=\sum_{k=1}^{d}\left[V_{d}^{\top}\right]_{i,k}\left[V_{d}\right]_{k,j}=\sum_{k=1}^{d}\left[V_{d}\right]_{k,i}\left[V_{d}\right]_{k,j}==\sum_{k=1}^{d}\frac{2}{i}\sin(i\tau_{k})\sin(i\rho)\frac{2}{j}\sin(j\tau_{k})\sin(j\rho)$
$=\frac{4\sin(i\rho)\sin(j\rho)}{ij}\sum_{k=1}^{d}\sin(i\tau_{k})\sin(j\tau_{k}).$
By Werner’s formula, we rewrite any term in the sum as
$\sin(i\tau_{k})\sin(j\tau_{k})=\frac{1}{2}\left[\cos((i-j)\tau_{k})-\cos((i+j)\tau_{k})\right].$
Consider first the case $i=j$, i.e. the diagonal terms of $V_{d}^{\top}V_{d}$. One has
$\sum_{k=1}^{d}\sin(i\tau_{k})\sin(j\tau_{k})$ $=\frac{1}{2}\sum_{k=1}^{d}\left[\cos((i-i)\tau_{k})-\cos((i+i)\tau_{k})\right]=\frac{1}{2}\sum_{k=1}^{d}\left[\cos(0)-\cos(2i\tau_{k})\right]$
$=\frac{1}{2}\sum_{k=1}^{d}1-\frac{1}{2}\sum_{k=1}^{d}\cos(2i\tau_{k})=\frac{d}{2}-0=\frac{d}{2},$
where the equality $\sum_{k=1}^{d}\cos(2i\tau_{k})=0$ is an easy (and well-known) consequence of the symmetry of the Chebyshev nodes. The symmetry of the Chebyshev nodes further implies that, for $i\neq j$,
$\sum_{k=1}^{d}\cos\left((i-j)\tau_{k}\right)=\sum_{k=1}^{d}\cos\left((i+j)\tau_{k}\right)=0.$
Hence one has
$\left[V_{d}^{\top}V_{d}\right]_{i,j}=\left[\frac{4}{ij}\sin(i\rho)\sin(j\rho)\right]\frac{d}{2}\delta_{i,j}=\left[\frac{2d}{ij}\sin(i\rho)\sin(j\rho)\right]\delta_{i,j},$
which shows the claim.
Proposition 3.1 is the key ingredient to unveil some features of $V_{d}$ and the corresponding matrix sequence. The first easy consequence is an exact formula for the determinant of $V_{d}$. This formula assumes a peculiar form for $\rho=\pi/(2d)$. The relevance of this case is described in [4], and will give further simplifications also for the next sections.
@@
(11) $\det V_{d}=\sqrt{\frac{(2d)^{d}}{(d!)^{2}}\prod_{i=1}^{d}\sin^{2}\left(i\rho\right)}.$
In particular, if $\rho=\pi/(2d)$, one has
$\det V_{d}=\sqrt{\frac{d(2d)^{d}}{(d!2^{d-1})^{2}}}.$
Proof 3.4.
@@
$\det V_{d}=\sqrt{\det(V_{d}^{\top}V_{d})}=\sqrt{\prod_{i=1}^{d}[V_{d}^{\top}V_{d}]_{i,i}}=\sqrt{\prod_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)}.$
Let us examine the argument of the last square root. One has
(12) $\prod_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)=\frac{\prod_{i=1}^{d}2d}{\prod_{i=1}^{d}i^{2}}\prod_{i=1}^{d}\sin^{2}(i\rho)=\frac{(2d)^{d}}{(d!)^{2}}\prod_{i=1}^{d}\sin^{2}(i\rho),$
which gives the first claim. For what concerns the specialization to $\rho=\pi/(2d)$, we shall invoke the trigonometric equality
(13) $\prod_{i=1}^{2d-1}\sin\left(\frac{i\pi}{2d}\right)=\frac{2d}{2^{2d-1}}.$
@@
which immediately yields the claim by taking the square root.
We close this section with an interesting fact. The identity (13) is known, but the authors struggled to find a reference. In fact, it appears in Wolfram MathWorld as a “personal communication”. For completeness, we report a possible proof.
@@
Evaluating the expression at $z=1$ we get the claim.

### v2

@@ 2.3 Segments with uniform length: the class (C4)
The class (C4) shows that the way segments overlap is a rather weak information for the purpose of unisolvence. In fact, we are going to show that the class (C4) is unisolvent for histopolation in $\mathscr{P}_{d-1}(I)$ whenever the nodes $\xi_{i}$, $i=1,\ldots,d$, are pairwise distinct (i.e., unisolvent for nodal interpolation), with no restriction on the measure of $s$. This can be seen as a consequence of the following general fact.
@@
Let $p:\mathbb{R}\to\mathbb{R}$ be a polynomial, and let $s$ be an interval. Define the translated of $s$ as $s+\xi:=\{x+\xi:\ x\in s\}$. If
$\int_{s+\xi}p(x)\mathrm{d}x=0$
for each $\xi\in\mathbb{R}$, then $p(x)=0$ for each $x\in\mathbb{R}$.
Proof.
Let $P$ be any antiderivative of $p$. Obviously, since $p$ is a polynomial, so is $P$. Then, writing the reference segment as $s=[\alpha,\beta]$, one has
$0=\int_{s+\xi}p(x)\mathrm{d}x=\int_{\alpha+\xi}^{\beta+\xi}p(x)\mathrm{d}x=P(\beta+\xi)-P(\alpha+\xi),$
which implies that $P(\beta+\xi)=P(\alpha+\xi)$ for each $\xi\in\mathbb{R}$. As a consequence, $P$ is a periodic function of period $|s|=\beta-\alpha$, and since it is a polynomial, it must be the constant polynomial. Since $P$ is an antiderivative of $p$, one has $p=P^{\prime}$, and since $P$ is constant it follows that $p(x)=P^{\prime}(x)=0$ for each $x\in\mathbb{R}$. ∎
@@
Let $s$ be a segment, and define $s+\xi:=\{x+\xi:\ x\in s\}$. Consider $\left\{\xi_{1},\ldots,\xi_{d}\right\}$ pairwise distinct nodes and let $p_{d-1}\in\mathscr{P}_{d-1}(I)$. If
@@
By the change of variable, one sees that this is a polynomial $q_{d-1}(\xi)$ of degree d1d-1 in the variable $\xi$. By hypothesis, this polynomial vanishes on the grid $\left\{\xi_{1},\ldots,\xi_{d}\right\}$, which consists of distinct nodes. Hence the polynomial $q_{d-1}(\xi)$ vanishes on the whole real line $\mathbb{R}$, implying that
$\int_{s+\xi}p_{d-1}(x)\mathrm{d}x=0$
for each $\xi\in\mathbb{R}$. By Lemma 2.8, one gets $p_{d-1}(x)=0$ for each $x\in\mathbb{R}$. ∎
@@
Proposition 2.9 also has a different proof. To this end, observe that the set of zeros of $p_{d-1}(x)$ is a discrete set, containing at most d1d-1 nodes. If it had more, then $p_{d-1}$ would be the zero polynomial. Call $\xi_{\max}$ the largest one, i.e. the node $\xi_{i}$ such that $\xi_{i}\geq\xi_{j}$ for any $j=1,\ldots,d-1$. Then, the result of Lemma 2.8 can be deduced as there exists $\xi\in\mathbb{R}$ such that $\xi-|s|\geq\xi_{\max}$. Hence, the sign of $p_{d-1}$ is constant in the interior of $s+\xi$, and since $\int_{s+\xi}p_{d-1}(x)\mathrm{d}x=0$, one has $p_{d-1}(x)=0$ for each $x$ in the (non empty) interior of $s+\xi$. Since the only polynomial vanisihing in an open and non empty set is the zero polynomial, this implies that $p_{d-1}(x)=0$ for each $x\in\mathbb{R}$.
@@ 2.4 Extension to generic intervals
and $\varepsilon$ is chosen so that $1+\varepsilon<2$. ∎
@@
The picture given in Theorems 2.5 and 2.11 and in Remarks 2.1 and 2.12 is even more general. In fact, if we consider sets $\{i\cdot\alpha_{i}^{(d)}=1,\ldots,d\}$, $\{i\cdot\beta_{i}^{(d)}=1,\ldots,d\}$, of natural numbers, in which for $d_{1}\neq d_{2}$ we have in general $\alpha_{i}^{(d_{1})}\neq\alpha_{i}^{(d_{2})}$ and $\beta_{i}^{(d_{1})}\neq\beta_{i}^{(d_{2})}$, $i\leq\min\{d_{1},d_{2}\}$, all the proofs work unchanged. This represents a further degree of generality in the analysis.
@@ 2.5 Some remarks on the independence of $c$ on $d$
In the non-normalized case, we have assumed the existence of a “separation distance” $c>0$ such that $\max_{i}\left(\beta_{i}-\alpha_{i}\right)\geq c$. We asked such a constant to be independent of $d$. This is clearly motivated by the fact that, if $c$ depended on $d$, it could decay to zero fast enough to simplify the exponential term $2^{d-2}$. It is plain to observe that, if $c$ depends on $d$ and does not decay exponentially, the exponential trend of the conditioning of the monomial basis persists. An example is provided when considering equidistributed segments in I=[1,1]I=[-1,1].
Example 2.15.
Let $I$ be an interval, and let $s_{1},\ldots,s_{d}$ be equidistributed segments covering $I$. Clearly, one has $|s_{i}|=\beta_{i}-\alpha_{i}=|I|/d$ for each $i=1,\ldots,d$, and this quantity tends to zero linearly with $d$ as $d$ grows, contradicting the hypotheses on $c$ required in the statement of Theorem 2.5. However, the conditioning is still exponential.
Since
$\|V_{d}\|_{2}=\max_{\begin{subarray}{c}z\in\mathbb{R}^{d}\\ z\neq 0\end{subarray}}\frac{\|V_{d}z\|_{2}}{\|z\|_{2}},$
by taking $z=e_{1}$, the first element of the canonical basis of $\mathbb{R}^{d}$, one has
$\|V_{d}\|_{2}\geq\|V_{d}e_{1}\|_{2}=\sqrt{\sum_{i=1}^{d}(\beta_{i}-\alpha_{i})^{2}}=\sqrt{\sum_{i=1}^{d}\left(\frac{|I|}{d}\right)^{2}}=\sqrt{d\left(\frac{|I|}{d}\right)^{2}}=\frac{|I|}{\sqrt{d}}.$
Merging this with the estimate for $\|V_{d}^{-1}\|_{2}$ given in Theorem 2.5, which does not depend on the size of the segments, we get
$\kappa_{2}(V_{d})=\|V_{d}\|_{2}\|V_{d}^{-1}\|_{2}\geq\frac{|I|}{\sqrt{d}}\frac{2^{d-1}}{\sqrt{d}}=|I|\frac{2^{d-2}}{d}.$
Notice that this represents a limiting value for all collection of segments that are arranged into chains, and that this result does not require specific hypotheses on $I$.
Suppose, in complete contrast, that such a constant $c=\max_{i}(\beta_{i}-\alpha_{i})$ is equal to zero; this means that every segment $s_{i}$ collapses to a point $\xi_{i}$. Even if these nodes are distinct, the problem (1) becomes ill-posed, as the matrix (2) becomes the zero matrix. This situation can be handled by placing the normalization factor $|s_{i}|$ in the histopolation conditions and considering the normalized case. If one imposes that the limiting nodes $\{\xi_{1},\ldots,\xi_{d}\}$ are pairwise distinct, by an immediate continuity argument and the mean value theorem, one deduces that the problem has a unique solution [7], and Corollary 2.6 implies that the corresponding matrix is exponentially ill-conditioned. However, the same continuity argument shows that the Vandermonde matrix (2) converges (in an entry-wise sense) to the nodal Vandermonde matrix based upon the nodes $\{\xi_{1},\ldots,\xi_{d}\}$ and written with respect to the monomial basis. The well-established results concerning the ill-conditioning of the monomial basis for the nodal theory [3, 22] therefore give a different proof of the exponential ill-conditioning in this limit.
@@ 3 Well-conditioning: Chebyshev basis and segments
In stark contrast to Section 2, a very different situation appears when the Chebyshev polynomials of the second kind
$U_{j-1}(x)=\frac{\sin(j\arccos x)}{\sin(\arccos x)}$
@@
is usually adopted. The two families are related via differentiation, as $\mathrm{d}/\mathrm{d}x\left(T_{j}(x)\right)=jU_{j-1}(x)$. As computed in [6, Section $2.2$], the Vandermonde matrix (2) assumes a very peculiar form with respect to the basis $U_{j-1}(x)$:
$\left[V_{d}\right]_{i,j}$ $=\int_{\alpha_{i}}^{\beta_{i}}U_{j-1}(x)\mathrm{d}x=\frac{1}{j}\int_{\alpha_{i}}^{\beta_{i}}\frac{\mathrm{d}T_{j}(x)}{\mathrm{d}x}\mathrm{d}x=\frac{1}{j}\left(T_{j}(\beta_{i})-T_{j}(\alpha_{i})\right)$ (6)
@@
$s_{i}=[\alpha_{i},\beta_{i}]=[\cos(\tau_{i}+\rho),\cos(\tau_{i}-\rho)],$ (7)
being $\tau_{i}$ the (pre-image of the) $i$-th Chebyshev node $\tau_{i}=\frac{(i-1/2)\pi}{d}$ for $i=1,\ldots,d$ and $\rho>0$ an arc length veryfing the existence conditions established in [6, Section 3.2]. Plugging the definition of the polynomial $U_{j-1}(x)$ and class (C2) segments into (6), one immediately obtains that the respective Vandermonde matrix is
$\left[V_{d}\right]_{i,j}=\frac{2}{j}\sin(j\tau_{i})\sin(j\rho).$ (8)
@@
$\left[V_{d}^{\top}V_{d}\right]_{i,i}=\begin{cases}\frac{2d}{i^{2}}\sin^{2}(i\rho)&\text{if }i=1,\ldots,d-1,\\ \frac{4d}{i^{2}}\sin^{2}(i\rho)&\text{if }i=d.\end{cases}$ (9)
Notice that, when $i=d$, the diagonal term is $4d/i^{2}\sin^{2}(i\rho)=4/d\sin^{2}(d\rho)$.
Proof.
@@
$\left[V_{d}^{\top}V_{d}\right]_{i,j}$ $=\sum_{k=1}^{d}\left[V_{d}^{\top}\right]_{i,k}\left[V_{d}\right]_{k,j}=\sum_{k=1}^{d}\left[V_{d}\right]_{k,i}\left[V_{d}\right]_{k,j}$ (10)
$=\sum_{k=1}^{d}\frac{2}{i}\sin(i\tau_{k})\sin(i\rho)\frac{2}{j}\sin(j\tau_{k})\sin(j\rho)=\frac{4\sin(i\rho)\sin(j\rho)}{ij}\sum_{k=1}^{d}\sin(i\tau_{k})\sin(j\tau_{k}).$
To conclude the proof, we shall establish the discrete orthogonality
$\sum_{k=1}^{d}\sin(l\tau_{k})\sin(m\tau_{k})=\begin{cases}\frac{d}{2}&\text{if }l=m<d\\ d&\text{if }l=m=d\\ 0&\text{if }l\neq m.\end{cases}$ (11)
To exploit the features of the Chebyshev nodes, we use Werner’s formula
$\sin(l\tau_{k})\sin(m\tau_{k})=\frac{1}{2}\left[\cos((l-m)\tau_{k})-\cos((l+m)\tau_{k})\right].$
Consider first the case $l=m$, i.e. the diagonal terms of $V_{d}^{\top}V_{d}$. One has
$\sum_{k=1}^{d}\sin(l\tau_{k})\sin(l\tau_{k})$ $=\frac{1}{2}\sum_{k=1}^{d}\left[\cos((l-l)\tau_{k})-\cos((l+l)\tau_{k})\right]=\frac{1}{2}\sum_{k=1}^{d}\left[\cos(0)-\cos(2l\tau_{k})\right]$
$=\frac{1}{2}\sum_{k=1}^{d}1-\frac{1}{2}\sum_{k=1}^{d}\cos(2l\tau_{k})=\frac{d}{2}-\frac{1}{2}\sum_{k=1}^{d}\cos(2l\tau_{k}).$ (12)
Recalling that $\tau_{k}=(2k-1)\pi/(2d)$, we compute the sum $\sum_{k=1}^{d}\cos(2l\tau_{k})$. Denoting by $\Re$ the real part of a complex number and by $\iota$ the imaginary unit, by Euler’s formula one has
$\sum_{k=1}^{d}\cos\left(\frac{l\pi}{d}(2k-1)\right)$ $=\sum_{k=1}^{d}\Re\left(e^{\iota\frac{l\pi}{d}(2k-1)}\right)=\Re\left(\sum_{k=1}^{d}e^{\iota\frac{l\pi}{d}(2k-1)}\right)$
$=\Re\left(e^{-\iota\frac{l\pi}{d}}\sum_{k=1}^{d}e^{\iota\frac{l\pi 2k}{d}}\right)=\Re\left(e^{-\iota\frac{l\pi}{d}}\sum_{k=1}^{d}\left(e^{\iota\frac{l2\pi}{d}}\right)^{k}\right).$
For $l=d$, one obtains $e^{i\frac{l2\pi}{d}}=e^{i\frac{d2\pi}{d}}=e^{i2\pi}=\cos(2\pi)=1$. Hence the real part can be removed:
$\sum_{k=1}^{d}e^{\iota\frac{d\pi}{d}(2k-1)}=e^{-\iota\pi}\sum_{k=1}^{d}1^{k}=-d.$
Plugging this into (3) one obtains the claim for $l=m=d$.
The case $1\leq l<d$ can be treated as a geometric series of ratio $e^{\iota\frac{l2\pi}{d}}$, which gives
$e^{-\iota\frac{l\pi}{d}}\sum_{k=1}^{d}\left(e^{\iota\frac{l2\pi}{d}}\right)^{k}=e^{-\iota\frac{l\pi}{d}}e^{\iota\frac{l2\pi}{d}}\frac{1-\left(e^{\iota\frac{l2\pi}{d}}\right)^{d}}{1-e^{\iota\frac{l2\pi}{d}}}=e^{-\iota\frac{l\pi}{d}}e^{\iota\frac{l2\pi}{d}}\frac{1-e^{\iota l2\pi}}{1-e^{\iota\frac{l2\pi}{d}}}.$ (13)
Since $1\leq l<d$, $2\pi l$ is an integer multiple of $2\pi$, while $2\pi l/d$ is not, and thus $e^{i\frac{l2\pi}{d}}\neq 1$ for any $l=1,\ldots,d-1$. It follows that
$\frac{1-e^{\iota l2\pi}}{1-e^{\iota\frac{l2\pi}{d}}}=\frac{1-1}{1-e^{\iota\frac{l2\pi}{d}}}=0.$
Plugging this into (3) we retrieve the first equality of (11). Therefore, the claim on the diagonal terms of $V_{d}^{\top}V_{d}$ is established.
We are left to prove that the off-diagonal terms of $V_{d}^{\top}V_{d}$ vanish. We hence show that
$\sum_{k=1}^{d}\cos\left((l-m)\tau_{k}\right)-\sum_{k=1}^{d}\cos\left((l+m)\tau_{k}\right)=0.$
To prove the above equality, we shall use the trigonometric representation of complex numbers again. Owing to the linearity of the real part, we focus on
$\sum_{k=1}^{d}\cos\left(\frac{n(2k-1)\pi}{d}\right)=\Re\left(\sum_{k=1}^{d}e^{\iota\frac{n}{2d}(2k-1)\pi}\right),$
either for n=lmn=l-m or $n=l+m$. We are going to show that both sums are equal to zero (when both lml-m and $l+m$ are even), or that both sums add to a completely imaginary number (when both lml-m and $l+m$ are odd). This reasoning clearly exhausts all cases.
Following the same reasoning that lead to (13), one finds
$\sum_{k=1}^{d}e^{\iota\frac{n}{2d}(2k-1)\pi}=e^{-\iota\frac{n\pi}{2d}}\sum_{k=1}^{d}\left(e^{\iota\frac{n\pi}{d}}\right)^{k}=e^{-\iota\frac{n\pi}{2d}}e^{\iota\frac{n\pi}{d}}\frac{1-e^{\iota n\pi}}{1-e^{\iota\frac{n\pi}{d}}}.$ (14)
The rightmost hand side of (14) exists as $n$ cannot be an entire multiple of $2d$ (because $l$ and $m$ represent off-diagonal terms, and so d<lm<d-d<l-m<d with $l-m\neq 0$ and $0<l+m<2d$), and this implies that $e^{\iota\frac{n\pi}{d}}\neq 1$.
To conclude the proof we consider separately the even and the odd case. First, suppose that n=lmn=l-m is even. In this case $e^{\iota n\pi}=\cos(2\pi)+\iota\sin(2\pi)=0$, and as a consequence
$\sum_{k=1}^{d}e^{\iota\frac{n}{2d}(2k-1)\pi}=e^{-\iota\frac{n\pi}{2d}}e^{\iota\frac{n\pi}{d}}\frac{1-e^{\iota n\pi}}{1-e^{\iota\frac{n\pi}{d}}}=e^{\iota\frac{n\pi}{2d}}\frac{1-1}{1-e^{\iota\frac{n\pi}{d}}}=0.$
Assume then n=lmn=l-m is odd. In such a case, $e^{\iota n\pi}=\cos(\pi)+\iota\sin(\pi)=-1$. Treating the sum as in (14) and manipulating, one gets
$\sum_{k=1}^{d}e^{\iota\frac{n}{2d}(2k-1)\pi}$ $=e^{-\iota\frac{n\pi}{2d}}e^{\iota\frac{n\pi}{d}}\frac{1-e^{\iota n\pi}}{1-e^{\iota\frac{n\pi}{d}}}=e^{\iota\frac{n\pi}{2d}}\frac{1+1}{1-e^{\iota\frac{n\pi}{d}}}=2\frac{e^{\iota n\frac{\pi}{2d}}}{1-e^{\iota n\frac{\pi}{d}}}=\frac{2}{e^{-\iota n\frac{\pi}{2d}}-e^{\iota n\frac{\pi}{2d}}}.$
Substituting the identity $e^{\iota n\frac{\pi}{2d}}-e^{-\iota n\frac{\pi}{2d}}=2\iota\sin(\frac{n\pi}{2d})$ and simplifying, one concludes that
$\sum_{k=1}^{d}e^{\iota\frac{n}{2d}(2k-1)\pi}=\frac{\iota}{\sin(\frac{n\pi}{2d})},$
which has no real part. The same argument applies to $\sum_{k=1}^{d}\cos\left((l+m)\tau_{k}\right)$, whence it follows that
$\Re\left(\sum_{k=1}^{d}\cos\left((l-m)\tau_{k}\right)\right)=\Re\left(\sum_{k=1}^{d}\cos\left((l+m)\tau_{k}\right)\right)=0.$
This immediately implies the last equality of (11). Gathering all the computations and plugging them into (10), one obtains
$\left[V_{d}^{\top}V_{d}\right]_{i,j}=\begin{cases}\frac{2d}{i^{2}}\sin^{2}(i\rho)&\text{if }i=1,\ldots,d-1,\\ \frac{4d}{i^{2}}\sin^{2}(i\rho)&\text{if }i=d,\\ 0&\text{if }i\neq j,\end{cases}$
which shows the claim. ∎
Proposition 3.1 is the key ingredient to unveil some features of $V_{d}$ and the corresponding matrix sequence. The first easy consequence is an exact formula for the modulus of the determinant of $V_{d}$, when the class (C2) is considered in combination with Chebyshev polynomials of the second kind. The sign of $\det V_{d}$ depends on $d$, and due to the reverse ordering of the Chebyshev nodes it equals $(-1)^{\frac{d(d-1)}{2}}$.
@@
$\left|\det V_{d}\right|=\sqrt{2\frac{(2d)^{d}}{(d!)^{2}}\prod_{i=1}^{d}\sin^{2}\left(i\rho\right)}.$ (15)
Proof.
@@
$\left|\det V_{d}\right|$ $=\sqrt{\det(V_{d}^{\top}V_{d})}=\sqrt{\prod_{i=1}^{d}[V_{d}^{\top}V_{d}]_{i,i}}=\sqrt{\left(\prod_{i=1}^{d-1}\frac{2d}{i^{2}}\sin^{2}(i\rho)\right)\frac{4d}{d^{2}}\sin^{2}(d\rho)}$
$=\sqrt{2\prod_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)}.$
Let us examine the argument of the last square root, up to the factor $2$. One has
$\prod_{i=1}^{d}\frac{2d}{i^{2}}\sin^{2}(i\rho)=\frac{\prod_{i=1}^{d}2d}{\prod_{i=1}^{d}i^{2}}\prod_{i=1}^{d}\sin^{2}(i\rho)=\frac{(2d)^{d}}{(d!)^{2}}\prod_{i=1}^{d}\sin^{2}(i\rho),$ (16)
which gives the claim. ∎
Formula (15) simplifies for $\rho=\pi/(2d)$. The relevance of this case, which lies at the intersection of classes (C1) and (C2), is described in [6]. This value will play an important role also in the next sections.
Corollary 3.3.
If $\rho=\pi/(2d)$, then (15) reduces to
$\left|\det V_{d}\right|=\sqrt{\frac{d^{d+1}}{(d!)^{2}2^{d-3}}}.$
Proof.
To prove this fact, we shall invoke the trigonometric equality
$\prod_{i=1}^{2d-1}\sin\left(\frac{i\pi}{2d}\right)=\frac{2d}{2^{2d-1}}.$ (17)
@@
which immediately yields the claim by including the factor $2$, simplifying and taking the square root.∎
We close this section with an interesting fact. The identity (17) is known, but the authors struggled to find a reference. For completeness, we report a possible proof.
@@
Evaluating the expression at $z=1$ we get the claim. ∎